10 parts
Longman Company manufactures shirts. During June, Longman made 1,100 shirts but had budgeted production at 1,225 shirts. Longman gath (Click on the icon to view the data.) Read the requirements. 13. Calculate the variable overhead cost variance. Select the formula, then enter the amounts and compute the cost variance for variable overhead (VOH) and identify whether the variance is fa VOH Cost Variance X Requi Data Table Calculate t 13. Varia 14. Varia $0.70 per DLHr Variable overhead cost standard 15. Total 4.50 DLHr per shirt Direct labor efficiency standard Actual amount of direct labor hours 5,030 DLHr Actual cost of variable overhead $4,024 $0.15 per DLHr Fixed overhead cost standard Budgeted fixed overhead $827 Actual cost of fixed overhead $852 Print Done Choose from any drop-down list and then click Check Answer. 10 parts remaining Clear All
ngman gathered the following additional data: ariance is favorable (F) or unfavorable (U) i Requirements X Calculate the following variances: 13. Variable overhead cost variance 16. Fixed overhead cost variance 14. Variable overhead efficiency variance 17. Fixed overhead volume variance 15. Total variable overhead variance 18. Total fixed overhead variance Print Done ar All Check Answer ook Pro
Answer #1

Longman company-variances VOH cost variance ( standard rate - actual rate) actual hours (0.7-0.8) 5030 { AR = 4024/ 5030 = 0.